# Clever: A Curated Benchmark for Formally Verified Code Generation

### v3

@@ 4. Finally, the correctness theorem statement in Lean 4:
{mdframed}
[roundcorner=10pt]
example_user
Goals to prove:
[GOALS]
[GOAL] 1
impl (-x) = impl x
(impl x = 0 x = 0) (0 x impl (-x) + impl x = 2 * x) (x 0 impl (-x) + impl x = -(2 * x))
[HYPOTHESES] 1
[HYPOTHESIS] impl :
[HYPOTHESIS] h_generated_spec : (x : ℤ), impl x = if x < 0 then -x else x
[HYPOTHESIS] x :
[HYPOTHESIS] h_x_lt_0 : x < 0
[HYPOTHESIS] h_not_0_lt_x : ¬0 < x
[HYPOTHESIS] h_impl : impl x = -x
[HYPOTHESIS] h_0_le_impl_x : 0 impl x
[GOAL] 2
0 impl x
impl (-x) = impl x
(impl x = 0 x = 0) (0 x impl (-x) + impl x = 2 * x) (x 0 impl (-x) + impl x = -(2 * x))
[HYPOTHESES] 1
[HYPOTHESIS] impl :
[HYPOTHESIS] h_generated_spec : (x : ℤ), impl x = if x < 0 then -x else x
[HYPOTHESIS] x :
[HYPOTHESIS] h_x_lt_0 : ¬x < 0
[STEPS]
[STEP] unfold problem_spec
[STEP] unfold generated_spec
[STEP] simp
[STEP] intro impl
[STEP] apply Iff.intro
[STEP] intro h_prob_spec
[STEP] have h_0_le_impl_x: 0 impl x := by
[STEP] simp [h_impl]
[STEP] linarith
[STEP] simp [h_0_le_impl_x]
[LAST STEP]
linarith [h_impl, h_0_le_impl_x, h_not_0_lt_x]
[ERROR MESSAGE]
linarith failed to find a contradiction
case pos
impl :
h_generated_spec : (x : ℤ), impl x = if x < 0 then -x else x
x :
h_x_lt_0 : x < 0
h_not_0_lt_x : ¬0 < x
h_impl : impl x = -x
h_0_le_impl_x : 0 impl x
False
failed
[END]
example_assistant
[RUN TACTIC]
have h_impl_neg_x := h_generated_spec (-x)
[END]
@@
{mdframed}
[roundcorner=10pt]
You are a good Lean 4 programmer. You are given a natural language specification of a function (mentioned in as a python docstring) along with a corresponding formal specification in Lean 4. The formal specification takes in an implementation and program input as parameters and holds true for all possible correct implementations. Your task is to generate a Lean 4 definition with a mentioned signature. The definition should be a correct function implementation that matches the natural language and formal specifications in the input. Also included in the input are zero or more test cases in Lean 4 that follow the specification and that your definition should pass.
The input usually follows the following format:
1. First we state the natural language specification of the function in a docstring format:
[NL DESCRIPTION]
def <function_name>(<input_type>) -> <output_type>
<NL Description>
2. Followed by the formal specification in Lean 4:
4. Finally, the test cases in Lean 4:
[TEST CASES]
#test implementation <input_1> = <expected_output_1>
#test implementation <input_2> = <expected_output_2>
You can first think about the problem in a general way and then write the definition. You can also use the following template to help you with the definition generation:
[THOUGHTS]
The definition should be a function that takes in an input
We can use a recursive helper function to ....
[END THOUGHTS]
Finally, write the generated implementation in the following format:
[GENERATED IMPLEMENTATION]
-- Change the following lines with actual generated formal implementation
let rec loop (<input_1>: <input_1_type>) (<input_2>: <input_2_type>) : <output_type> := ....
[END]
Please closely follow the format as shown in the examples below. Make sure that your response always ends with [END]. Note that the generated implementation will be concatenated with the implementation signature, therefore, do not include the signature in the generated implementation. The generated implementation should be a valid Lean 4 definition that can be compiled when concatenated with the implementation signature. Try to write an implementation where termination can be automatically verified; for example, always use library functions (‘Int.lcm‘, String.find‘, Nat.fermatNumber‘, etc.) or match statements if possible. DO NOT ever use the in keyword, it is not a valid keyword in Lean 4.
@@ 4. Finally, the test cases in Lean 4:
{mdframed}
[roundcorner=10pt]
`example_user`
[NL DESCRIPTION]
def find_magnitude(x: int) -> int
Given an integer x, your task is to find the magnitude of x.
The magnitude of an integer is defined as the absolute value of the integer.
[SPECIFICATION]
def problem_spec
-- function signature
(impl: Int Int)
-- inputs
(x: Int) :=
-- spec
let spec (result: Int) :=
(result 0)
(impl (-x) = result)
(result = 0 x = 0)
(0 x impl (-x) + result = 2 * x)
(x 0 impl (-x) + result = -2 * x)
-- program terminates
result, impl x = result
-- return value satisfies spec
spec result
[IMPLEMENTATION SIGNATURE]
def implementation (x: Int) : Int :=
[TEST CASES]
#test implementation (-1) = 1
#test implementation 0 = 0
#test implementation 5 = 5
#test implementation (-100) = 100
#test implementation 100 = 100
`example_assistant`
[THOUGHTS]
If x is less than 0 then return ‘-x‘, otherwise return x‘.
[END THOUGHTS]
[GENERATED IMPLEMENTATION]
if x < 0 then -x else x
[END]
@@
{mdframed}
[roundcorner=10pt]
You are a good Lean 4 programmer. You are given:
1. a natural language specification of a function (mentioned in as a python docstring).
2. a corresponding problem specification in lean 4.
3. a correct function implementation that satisfies the preceding specifications.
Your task is to write a formal proof in Lean 4 that the function implementation is correct and satisfies the formal specification.
The correctness statement is stated in the following format:
1. First we state the natural language description of the function in a docstring format:
[NL DESCRIPTION]
def <function_name>(<input_type>) -> <output_type>
<NL Description>
2. Followed by the formal specification in Lean 4:
4. Finally, the correctness theorem statement in Lean 4:
[THEOREM STATEMENT]
theorem correctness
(input: <input_type>)
: problem_spec implementation input
You can first think about the problem in a general way and then write the proof. You can also use the following template to help you with the proof generation:
[THOUGHTS]
The implementation should match the problem_spec in the first case because ....
The implementation should match the problem_spec in the second case because ....
[END THOUGHTS]
Finally, write a proof in Lean 4 that the implementation is correct and satisfies the specification. The proof should be in the following format:
[PROOF]
-- proof tactics
by
rw [...]
simp ...
-- more proof tactics
[END]
Please closely follow the format as shown in the examples below. Make sure that your response always ends with [END]. The proof must be in Lean 4 and should be valid. Make sure to always start your proof with ‘[PROOF]‘ followed by a by keyword. The proof should end with ‘[END]‘. DO NOT forget to write the PROOF section. The proof section must be there in your response. Please do NOT use sorry in your proof anywhere. The proof must be complete and valid.
@@ 4. Finally, the correctness theorem statement in Lean 4:
{mdframed}
[roundcorner=10pt]
`example_user`
[NL DESCRIPTION]
def find_magnitude(x: int) -> int
Given an integer x, your task is to find the magnitude of x.
The magnitude of an integer is defined as the absolute value of the integer.
[SPECIFICATION]
def problem_spec
[IMPLEMENTATION]
def implementation (x: Int) : Int :=
if x < 0 then -x else x
[THEOREM STATEMENT]
theorem correctness
(x: Int)
: problem_spec implementation x :=
`example_assistant`
[THOUGHTS]
We can start by unfolding the problem_spec and assigning the implementations output to a temporary variable result‘.
Early on, we will want to break the proof up into cases based on whether x is positive, negative, or zero.
Many of the intermediate goals can likely be proven using linarith‘.
[END THOUGHTS]
[PROOF]
by
unfold problem_spec
let result := implementation x
use result
simp [result]
simp [implementation]
by_cases h_0_eq_x: x = 0
-- if x = 0 then
simp [h_0_eq_x]
by_cases h_x_lt_0: x < 0
-- if x < 0 then
simp [h_x_lt_0]
apply And.intro
linarith
apply And.intro
intro; linarith
by_cases h_0_lt_x: 0 < x
-- if 0 < x then
simp [h_0_lt_x]
linarith
have h_x_eq_0: x = 0 := by
linarith
simp [h_x_eq_0]
intro h_x_le_0
simp at h_x_lt_0
have h_x_eq_0: x = 0 := by
linarith
contradiction
[END]
@@
{mdframed}
[roundcorner=10pt]
`example_user`
Goals to prove:
[GOALS]
[GOAL] 1
(0 if x < 0 then -x else x)
((if 0 < x then x else -x) = if x < 0 then -x else x)
((if x < 0 then -x else x) = 0 x = 0)
(0 x ((if 0 < x then x else -x) + if x < 0 then -x else x) = 2 * x)
(x 0 ((if 0 < x then x else -x) + if x < 0 then -x else x) = -(2 * x))
[HYPOTHESES] 1
[HYPOTHESIS] x :
[HYPOTHESIS] result : := implementation x
[STEPS]
[STEP] unfold problem_spec
[STEP] let result := implementation x
[STEP] use result
[STEP] simp [result]
[STEP] simp [implementation]
[INCORRECT STEPS]
[STEP] cases x with n
[LAST STEP]
cases lt_or_ge x 0 with hx hx
[ERROR MESSAGE]
unknown tactic
[END]
`example_assistant`
[RUN TACTIC]
by_cases h_0_eq_x: x = 0
[END]
@@ A.5 Some Proof Found
Figure 21 shows an example of a proof found for implementation certification by Claude-3.7 using COPRA.
{mdframed}
[roundcorner=6pt, innerleftmargin=6pt, innerrightmargin=6pt, innertopmargin=4pt, innerbottommargin=4pt]
def problem_spec (impl : Nat Nat) (n : Nat) :=
let spec (result : Nat) :=
let factorial := Nat.factorial n;
(0 < n result / factorial = impl (n - 1))
(n = 0 result = 1);
result, impl n = result spec result
def implementation (n : Nat) : Nat :=
match n with
| 0 => 1
| n+1 => Nat.factorial (n+1) * implementation n
#test implementation 4 = 288
theorem correctness (n : Nat) : problem_spec implementation n := by
unfold problem_spec
let result := implementation n
use result
simp [result]
apply And.intro
unfold implementation
simp [Nat.factorial_succ]
intro h_n_pos
cases n
exfalso
exact h_n_pos.false
rename_i n
simp [Nat.succ_eq_add_one, Nat.add_sub_cancel]
rw [Nat.factorial_succ]
have h_fac_pos : 0 < (n + 1) * n.factorial := by
exact Nat.mul_pos (Nat.succ_pos n) (Nat.factorial_pos n)
have h_cancel : (n + 1) * n.factorial * implementation n
/ ((n + 1) * n.factorial) = implementation n := by
rw [Nat.mul_div_cancel_left (implementation n) h_fac_pos]
simp [h_cancel]
unfold implementation
cases n
simp [Nat.factorial_zero]
rename_i n
simp [Nat.add_zero]
simp [Nat.factorial_succ]
left
rw [implementation.eq_def]
simp [Nat.mul_assoc]
cases n
simp [Nat.zero_eq]
rename_i n
simp [Nat.factorial_succ]
rw [Nat.mul_assoc]
intro h_n_eq_0
rw [h_n_eq_0, implementation]

### v4

@@ A.4 Baseline Prompts
example_user Goals to prove: [GOALS] [GOAL] 1 impl (-x) = impl x (impl x = 0 x = 0) (0 x impl (-x) + impl x = 2 * x) (x 0 impl (-x) + impl x = -(2 * x)) [HYPOTHESES] 1 [HYPOTHESIS] impl : [HYPOTHESIS] h_generated_spec : (x : ℤ), impl x = if x < 0 then -x else x [HYPOTHESIS] x : [HYPOTHESIS] h_x_lt_0 : x < 0 [HYPOTHESIS] h_not_0_lt_x : ¬0 < x [HYPOTHESIS] h_impl : impl x = -x [HYPOTHESIS] h_0_le_impl_x : 0 impl x [GOAL] 2 0 impl x impl (-x) = impl x (impl x = 0 x = 0) (0 x impl (-x) + impl x = 2 * x) (x 0 impl (-x) + impl x = -(2 * x)) [HYPOTHESES] 1 [HYPOTHESIS] impl : [HYPOTHESIS] h_generated_spec : (x : ℤ), impl x = if x < 0 then -x else x [HYPOTHESIS] x : [HYPOTHESIS] h_x_lt_0 : ¬x < 0 [STEPS] [STEP] unfold problem_spec [STEP] unfold generated_spec [STEP] simp [STEP] intro impl [STEP] apply Iff.intro [STEP] intro h_prob_spec ........ [STEP] have h_0_le_impl_x: 0 impl x := by [STEP] simp [h_impl] [STEP] linarith [STEP] simp [h_0_le_impl_x] [LAST STEP] linarith [h_impl, h_0_le_impl_x, h_not_0_lt_x] [ERROR MESSAGE] linarith failed to find a contradiction case pos impl : h_generated_spec : (x : ℤ), impl x = if x < 0 then -x else x x : h_x_lt_0 : x < 0 h_not_0_lt_x : ¬0 < x h_impl : impl x = -x h_0_le_impl_x : 0 impl x False failed [END] example_assistant [RUN TACTIC] have h_impl_neg_x := h_generated_spec (-x) [END]
@@
You are a good Lean 4 programmer. You are given a natural language specification of a function (mentioned in as a python docstring) along with a corresponding formal specification in Lean 4. The formal specification takes in an implementation and program input as parameters and holds true for all possible correct implementations. Your task is to generate a Lean 4 definition with a mentioned signature. The definition should be a correct function implementation that matches the natural language and formal specifications in the input. Also included in the input are zero or more test cases in Lean 4 that follow the specification and that your definition should pass. The input usually follows the following format: 1. First we state the natural language specification of the function in a docstring format: ‘‘‘ [NL DESCRIPTION] def <function_name>(<input_type>) -> <output_type> """ <NL Description> """ ‘‘‘ 2. Followed by the formal specification in Lean 4: ........ 4. Finally, the test cases in Lean 4: ‘‘‘ [TEST CASES] #test implementation <input_1> = <expected_output_1> #test implementation <input_2> = <expected_output_2> ‘‘‘ You can first think about the problem in a general way and then write the definition. You can also use the following template to help you with the definition generation: ‘‘‘ [THOUGHTS] The definition should be a function that takes in an input We can use a recursive helper function to .... [END THOUGHTS] ‘‘‘ Finally, write the generated implementation in the following format: ‘‘‘ [GENERATED IMPLEMENTATION] -- Change the following lines with actual generated formal implementation let rec loop (<input_1>: <input_1_type>) (<input_2>: <input_2_type>) : <output_type> := .... [END] ‘‘‘ Please closely follow the format as shown in the examples below. Make sure that your response always ends with [END]. Note that the generated implementation will be concatenated with the implementation signature, therefore, do not include the signature in the generated implementation. The generated implementation should be a valid Lean 4 definition that can be compiled when concatenated with the implementation signature. Try to write an implementation where termination can be automatically verified; for example, always use library functions (‘Int.lcm‘, String.find‘, Nat.fermatNumber‘, etc.) or match statements if possible. DO NOT ever use the in keyword, it is not a valid keyword in Lean 4.
@@
`example_user` [NL DESCRIPTION] def find_magnitude(x: int) -> int """ Given an integer x, your task is to find the magnitude of x. The magnitude of an integer is defined as the absolute value of the integer. """ [SPECIFICATION] def problem_spec -- function signature (impl: Int Int) -- inputs (x: Int) := -- spec let spec (result: Int) := (result 0) (impl (-x) = result) (result = 0 x = 0) (0 x impl (-x) + result = 2 * x) (x 0 impl (-x) + result = -2 * x) -- program terminates result, impl x = result -- return value satisfies spec spec result [IMPLEMENTATION SIGNATURE] def implementation (x: Int) : Int := [TEST CASES] #test implementation (-1) = 1 #test implementation 0 = 0 #test implementation 5 = 5 #test implementation (-100) = 100 #test implementation 100 = 100 `example_assistant` [THOUGHTS] If x is less than 0 then return ‘-x‘, otherwise return x‘. [END THOUGHTS] [GENERATED IMPLEMENTATION] if x < 0 then -x else x [END]
@@
You are a good Lean 4 programmer. You are given: 1. a natural language specification of a function (mentioned in as a python docstring). 2. a corresponding problem specification in lean 4. 3. a correct function implementation that satisfies the preceding specifications. Your task is to write a formal proof in Lean 4 that the function implementation is correct and satisfies the formal specification. The correctness statement is stated in the following format: 1. First we state the natural language description of the function in a docstring format: ‘‘‘ [NL DESCRIPTION] def <function_name>(<input_type>) -> <output_type> """ <NL Description> """ ‘‘‘ 2. Followed by the formal specification in Lean 4: ........ 4. Finally, the correctness theorem statement in Lean 4: ‘‘‘ [THEOREM STATEMENT] theorem correctness (input: <input_type>) : problem_spec implementation input := ‘‘‘ You can first think about the problem in a general way and then write the proof. You can also use the following template to help you with the proof generation: ‘‘‘ [THOUGHTS] The implementation should match the problem_spec in the first case because .... The implementation should match the problem_spec in the second case because .... .... [END THOUGHTS] ‘‘‘ Finally, write a proof in Lean 4 that the implementation is correct and satisfies the specification. The proof should be in the following format: ‘‘‘ [PROOF] -- proof tactics by rw [...] simp ... -- more proof tactics [END] ‘‘‘ Please closely follow the format as shown in the examples below. Make sure that your response always ends with [END]. The proof must be in Lean 4 and should be valid. Make sure to always start your proof with ‘[PROOF]‘ followed by a by keyword. The proof should end with ‘[END]‘. DO NOT forget to write the PROOF section. The proof section must be there in your response. Please do NOT use sorry in your proof anywhere. The proof must be complete and valid.
@@
`example_user` [NL DESCRIPTION] def find_magnitude(x: int) -> int """ Given an integer x, your task is to find the magnitude of x. The magnitude of an integer is defined as the absolute value of the integer. """ [SPECIFICATION] def problem_spec ........ [IMPLEMENTATION] def implementation (x: Int) : Int := if x < 0 then -x else x [THEOREM STATEMENT] theorem correctness (x: Int) : problem_spec implementation x := `example_assistant` [THOUGHTS] We can start by unfolding the problem_spec and assigning the implementations output to a temporary variable result‘. Early on, we will want to break the proof up into cases based on whether x is positive, negative, or zero. Many of the intermediate goals can likely be proven using linarith‘. [END THOUGHTS] [PROOF] by unfold problem_spec let result := implementation x use result simp [result] simp [implementation] by_cases h_0_eq_x: x = 0 -- if x = 0 then simp [h_0_eq_x] by_cases h_x_lt_0: x < 0 -- if x < 0 then simp [h_x_lt_0] apply And.intro linarith apply And.intro intro; linarith ........ by_cases h_0_lt_x: 0 < x -- if 0 < x then simp [h_0_lt_x] linarith have h_x_eq_0: x = 0 := by linarith simp [h_x_eq_0] intro h_x_le_0 simp at h_x_lt_0 have h_x_eq_0: x = 0 := by linarith contradiction [END]
@@
`example_user` Goals to prove: [GOALS] [GOAL] 1 (0 if x < 0 then -x else x) ((if 0 < x then x else -x) = if x < 0 then -x else x) ((if x < 0 then -x else x) = 0 x = 0) (0 x ((if 0 < x then x else -x) + if x < 0 then -x else x) = 2 * x) (x 0 ((if 0 < x then x else -x) + if x < 0 then -x else x) = -(2 * x)) [HYPOTHESES] 1 [HYPOTHESIS] x : [HYPOTHESIS] result : := implementation x [STEPS] [STEP] unfold problem_spec [STEP] let result := implementation x [STEP] use result [STEP] simp [result] [STEP] simp [implementation] [INCORRECT STEPS] [STEP] cases x with n [LAST STEP] cases lt_or_ge x 0 with hx hx [ERROR MESSAGE] unknown tactic [END] `example_assistant` [RUN TACTIC] by_cases h_0_eq_x: x = 0 [END]
@@ A.5 Some Proof Found
Figure˜22 shows an example of a proof found for implementation certification by Claude-3.7 using COPRA.
(a) def problem_spec (impl : Nat Nat) (n : Nat) := let spec (result : Nat) := let factorial := Nat.factorial n; (0 < n result / factorial = impl (n - 1)) (n = 0 result = 1); result, impl n = result spec result (b) def implementation (n : Nat) : Nat := match n with | 0 => 1 | n+1 => Nat.factorial (n+1) * implementation n (c) #test implementation 4 = 288 (d) theorem correctness (n : Nat) : problem_spec implementation n := by unfold problem_spec let result := implementation n use result simp [result] apply And.intro unfold implementation simp [Nat.factorial_succ] intro h_n_pos cases n exfalso exact h_n_pos.false rename_i n simp [Nat.succ_eq_add_one, Nat.add_sub_cancel] rw [Nat.factorial_succ] have h_fac_pos : 0 < (n + 1) * n.factorial := by exact Nat.mul_pos (Nat.succ_pos n) (Nat.factorial_pos n) have h_cancel : (n + 1) * n.factorial * implementation n / ((n + 1) * n.factorial) = implementation n := by rw [Nat.mul_div_cancel_left (implementation n) h_fac_pos] simp [h_cancel] unfold implementation cases n simp [Nat.factorial_zero] rename_i n simp [Nat.add_zero] simp [Nat.factorial_succ] left rw [implementation.eq_def] simp [Nat.mul_assoc] cases n simp [Nat.zero_eq] rename_i n simp [Nat.factorial_succ] rw [Nat.mul_assoc] intro h_n_eq_0 rw [h_n_eq_0, implementation]